Q = 40-0.05P Derive inverse demand What is demand choke price

  1. Q = 40-0.05P
  1. Derive inverse demand
  1. What is demand choke price
Answer #1

Answer
a)
inverse demand function is
Q=40-0.05P
0.05P=40-Q
P=800-20Q
the inverse demand function is P=800-20Q

b)
a choke price where Qd=0
P=800-20*0
P=800
the price is $800